OpenStudy (anonymous):

Multiple choice please help :) What is the magnitude and direction of Vector GH if G(3, -2) and H(6, 4)? magnitude: 6.71 units; direction 63.43° magnitude: 6.71 units: direction: 26.57° magnitude: 3.61 units; direction: 26.57° magnitude: 3.61 units: direction: 63.43° I know that its not the third one!

OpenStudy (ybarrap):

\[g = 3i - 2j\] \[h = 6i + 4j\] \[h-g = 3i + 6j\] (vector from g to h) \[\sqrt{3^{2}+9^{2}}\] (it's magnitude) \[\theta = \tan^{-1} 6/3\] (it's angle)

OpenStudy (anonymous):

hmmm im not sure which part i need to solve?

OpenStudy (ybarrap):

you'll just need to compute the square root of 3^2 + 9^2 and the arc tangent of 6/3

OpenStudy (ybarrap):

The first will be magnitude, the second will be the angle. Need a picture?

OpenStudy (anonymous):

the square root of 90 is 9.48 no what?

OpenStudy (anonymous):

now

OpenStudy (ybarrap):

sorry, I meant you'll just need to compute the square root of 3^2 + 6^2 and the arc tangent of 6/3. (I changed 9 to 6 here)

OpenStudy (ybarrap):

everything else looks good

OpenStudy (anonymous):

so 6.7 but i dont know how to calculate the tangent could you helo?

OpenStudy (ybarrap):

you should get \[\sqrt{3^{3}+6^{2}}=6.71\]

OpenStudy (anonymous):

im just confused on what to do next :/

OpenStudy (ybarrap):

For the magnitude

OpenStudy (ybarrap):

Do you have a picture in your head about what is required? It helps alot

OpenStudy (anonymous):

im lost lol

OpenStudy (ybarrap):

just a sec, I'll try to draw this

OpenStudy (fibonaccichick666):

may I interject, LaurenAshley, if you graph the vectors it usually helps you to understand. =)

OpenStudy (ybarrap):